You are given a binomial random variable with n = 25 and p = 0.35.
A. Calculate the mean for the random variable.
B. Calculate the standard deviation for the random variable.

Answer by stanbon(75887)

You are given a binomial random variable with n = 25 and p = 0.35.
A. Calculate the mean for the random variable:: u = np = 25*0.35 = 8.75
-------------------------
B. Calculate the standard deviation for the random variable::
std = sqrt(npq) = sqrt(8.75*0.65) = 2.3848
============================================
Cheers,
Stan H.